# Hodge diamonds of the Landau–Ginzburg orbifolds

### 26 Feb 2024 v2

@@
Abstract.
Consider the pairs $(f,G)$ with $f=f(x_{1},\dots,x_{N})$ being a polynomial defining a quasihomogeneous singularity and $G$ being a subgroup of ${\mathrm{SL}}(N,{\mathbb{C}})$, preserving $f$. In particular, $G$ is not necessary abelian. Assume further that $G$ contains the grading operator $j_{f}$ and $f$ satisfies the Calabi-Yau condition. We prove that the nonvanishing bigraded pieces of the B–model state space of $(f,G)$ form a diamond. We identify its topmost, bottommost, leftmost and rightmost entries as one-dimensional and show that this diamond enjoys the essential horizontal and vertical isomorphisms.
1. Introduction
Let a polynomial $f\in{\mathbb{C}}[x_{1},\dots,x_{N}]$ be quasihomogeneous with respect to some positive integers $d_{0},d_{1},\dots,d_{N}$, i.e.
$f(\lambda^{d_{1}}x_{1},\dots,\lambda^{d_{N}}x_{N})=\lambda^{d_{0}}f(x_{1},\dots,x_{N}),\quad\forall\lambda\in{\mathbb{C}}^{\ast}.$
Assume also that $x_{1}=\cdots=x_{N}=0$ is the only critical point of $f$ and $d_{1},\dots,d_{N}$ have no common factor. Then the zero set ${f(x_{1},\dots,x_{N})=0}$ defines a degree $d_{0}$ quasismooth hypersurface $X_{f}$ in ${\mathbb{P}}(d_{1},\dots,d_{N})$. Such hypersurfaces became of great interest in the early 90s in the context of mirror symmetry (cf. [BH93, BH95]). In particular, if the Calabi–Yau condition $d_{0}=\sum_{k=1}^{N}d_{k}$ holds, then the first Chern class of $X_{f}$ vanishes and, hence, its canonical bundle is trivial meaning that $X_{f}$ is a Calabi–Yau variety.
@@ 1. Introduction
1.1. Hodge diamonds of Calabi-Yau manifolds
The state space of a Calabi–Yau manifolds $X$ is the cohomology ring $H^{\ast}(X)$. This cohomology ring is naturally bigraded building up a Hodge diamond of size $D:=\dim_{\mathbb{C}}X$. In particular the following properties hold
@@ 1.1. Hodge diamonds of Calabi-Yau manifolds
    there is a “vertical” vector space isomorphism $H^{p,q}(X)\cong(H^{D-q,D-p}(X))^{\vee}$, where $(-)^{\vee}$ stands for the dual vector space
In physics, this bigrading is coming from considering Calabi-Yau B-model associated to $X$ and A-model bigrading is obtained from it by the so-called rotation of the diamond by $90^{\circ}$. On the level of state spaces the mirror symmetry map is an isomorphism of the cohomology vector spaces for a dual pair of Calabi-Yau manifolds switching A- and B- model bigradings.
More, generally if considering Calabi-Yau orbifolds in place of quasismooth varieties, one replaces the ordinary cohomology ring by the Chen–Ruan cohomology ring $H^{\ast}_{orb}$. This is an essential question if $H^{\ast}_{orb}$ forms a Hodge diamond too. Some of the Hodge diamond properties above follow directly from the definitions or could be verified in the similar way to our main Theorem 1 below, while the others (like the property (4)) were not investigated in literature up to our knowledge and do not look to be straightforward.
1.2. Landau–Ginzburg orbifolds
Another facet of mirror symmetry is given by matching the so-called Landau-Ginzburg orbifolds in place of Calabi-Yau manifolds or orbifolds (cf. [W93, IV90, Kreu94, V89]). Mathematically, these are the pairs $(f,G)$ with $f$ being a quasihomogeneous polynomial with the only critical point $0\in{\mathbb{C}}^{N}$ and $G$ being a group of symmetries of $f$.
@@ 1.2. Landau–Ginzburg orbifolds
${\mathrm{GL}}_{f}:=\left\{g\in{\mathrm{GL}}(N,{\mathbb{C}})\ |\ f(g\cdot{\bf x})=f({\bf x})\right\}.$
@@
$j_{f}\cdot(x_{1},\dots,x_{N}):=(e^{2\pi\sqrt{-1}d_{1}/d_{0}}x_{1},\dots,e^{2\pi\sqrt{-1}d_{N}/d_{0}}x_{N}).$
@@
For any $G\subseteq{\mathrm{GL}}_{f}$, the pair $(f,G)$ is called a Landau–Ginzburg orbifold. One associates to it the state space, which is the $G$-equivariant generalization of a Jacobian ring of $f$, together with A- and B- model bigradings, which again differ by $90^{\circ}$-rotation from one another provided $G$ acts by transfromations with determinant 1. Since these are interdependent, within this paper we will focus on just the B-model, as it has clearer geometric interpretation, much alike bigrading on the cohomology of Calabi-Yau manifold. For this reason we will call this state space endowed with B-model bigrading by ${\mathcal{B}}(f,G)$.
Up till now Landau–Ginzburg orbifolds were mostly investigated for the groups $G$ acting diagonally on ${\mathbb{C}}^{N}$ and also for $f$ belonging to a very special class of polynomials — the so–called invertible polynomials (cf. [K09, FJJS, BT2, K03, K06, BTW16, BTW17]). Also some work was done for the symmetry groups $G=S\ltimes G^{d}$ with $G^{d}$ acting diagonally and $S\subseteq S_{N}$ – a subgroup of a symmetric group on $N$ elements ([EGZ18, EGZ20, MO70, WWP, Y16, CJMPW23, BI21, BI22, I23]). We relax both conditions in this paper.
Once again, the mirror symmetry attempts to match A- and B- state spaces for dual pairs of Landau-Ginzburg orbifolds. It should be mentioned, that the state space enjoys several other structures besides being just bigraded vector space, like multiplication or bilinear form on both A- and B- sides, which should also be compatible with the mirror map. They will not be considered in the present paper.
1.3. Calabi–Yau/Landau–Ginzburg correspondence
The final piece of matching comes from Calabi–Yau/Landau–Ginzburg correspondence, which relates respective A-models and their state spaces coming from Calabi–Yau and Landau–Ginzburg geometries of $(X_{f},G/J)$ and $(f,G)$ provided $J\subseteq G$. Mathematically, up till now this was proved for diagonal symmetry groups by Chiodo and Ruan [CR11, Theorem 14] and in some special cases with $N=5$ in [M].
@@ 1.3. Calabi–Yau/Landau–Ginzburg correspondence
Theorem 1.
Let $f\in{\mathbb{C}}[x_{1},\dots,x_{N}]$ be a quasihomogeneous polynomial satisfying Calabi–Yau condition and defining an isolated singularity. Then for any $G\subseteq{\mathrm{SL}}_{f}$, such that $J\subseteq G$ the state space ${\mathcal{B}}(f,G)$ forms a diamond of size $N-2$ in a sense that it satisfies conditions (1)–(6) above.
@@
The proof is summed up in Propositions 35, 36 and 37. The vertical and horizontal isomorphism are given in Section 4. ∎
1.4. Acknowledgements
2. Preliminaries and notations
2.1. Quasihomogeneous singularities
@@
$f(\lambda^{d_{1}}x_{1},\dots,\lambda^{d_{N}}x_{N})=\lambda^{d_{0}}f(x_{1},\dots,x_{N}),\quad\forall\lambda\in{\mathbb{C}}.$ (QH1)
@@
We will say that $f$ defines an isolated singularity at $0\in{\mathbb{C}}^{N}$ if 00 is the only critical point of $f$. According to K. Saito [S71, Satz 1.3] one may consider without changing the singularity only the quasihomogeneous polynomials, such that $20<q_{k}\leq 1/$ 2 for all $k=1,\dots,N$. Moreover, we may assume that all variable $x_{k}$ with $q_{k}=1/2$ enter $f$ only in monomial $x_{k}^{2}$, in particular, there are no monomials of the form $x_{i}x_{j}$ with $i\neq j$. Then the number of its monomials is not less than $N$, the number of variables.
Example 2.
@@
[…]
Using the word ’type’ we mean the certain structure of the monomials without specifying the exponents $a_{i}$.
It’s easy to see that for Fermat, chain and loop type polynomials, the reduced weights $q_{1},\dots,q_{N}$ are defined in a unique way. This is also true for any quasihomogeneous singularity (cf. [S71, Korollar 1.7]). We have
$q_{1}=\frac{1}{a_{1}}\quad$ Fermat type, (1)
$q_{i}=\sum_{j=i}^{N}\frac{(-1)^{j-i}}{a_{1}\cdots a_{j}}\quad$ chain type, (2)
$q_{i}=(-1)^{N-1}\frac{1-a_{i}+a_{i}\sum_{k=2}^{N-1}(-1)^{k}\prod_{l=2}^{k}a_{i-l+1}}{\prod_{k=1}^{N}a_{k}-(-1)^{N}}\quad$ loop type, (3)
@@
Example 3 (Section 13.1 of [AGV85]).
If $N\leq 2$, then all quasihomogeneous isolated singularities are given by the $\oplus$–sums of the Fermat, chain or loop type polynomials.
Example 4 (Section 13.2 of [AGV85]).
If $N=3$, then all quasihomogeneous isolated singularities are given by the following polynomials $f_{I}=x_{1}^{a_{1}}+x_{2}^{a_{2}}+x_{3}^{a_{3}}$, $f_{II}=x_{1}^{a_{1}}+x_{2}^{a_{2}}x_{1}+x_{3}^{a_{3}}$ , $f_{III}=x_{1}^{a_{1}}+x_{2}^{a_{2}}x_{1}+x_{3}^{a_{3}}x_{1}+\varepsilon x_{2}^{p}x_{3}^{q}$, $f_{IV}=x_{1}^{a_{1}}+x_{2}^{a_{2}}x_{3}+x_{3}^{a_{3}}x_{2}$ , $f_{V}=x_{1}^{a_{1}}+x_{2}^{a_{2}}x_{1}+x_{3}^{a_{3}}x_{2}$, $f_{VI}=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}x_{1}+x_{3}^{a_{3}}x_{1}+\varepsilon x_{2}^{p}x_{3}^{q}$, $f_{VII}=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}x_{3}+x_{3}^{a_{3}}x_{1}$ with some positive $a_{1},a_{2},a_{3}$. The numbers $a_{i}$ are arbitrary for $f_{I},f_{II},f_{IV},f_{V},f_{VII}$, however the polynomials $f_{III}$ and $f_{VI}$ are only quasihomogeneous if $\varepsilon\neq 0$ and some additional combinatorial condition on $a_{1},a_{2},a_{3}$ holds. In particular the least common divisor of $(a_{2},a_{3})$ should be divisible by $a_{1}-1$ for $f_{III}$ to exist. Allowed values of $\varepsilon$ depend quiet heavily on $a_{i}$. In particular $\varepsilon\not\in\{0,\sqrt{-1},-\sqrt{-1}\}$ for $f=x_{1}^{3}+x_{1}x_{2}^{2}+x_{1}x_{3}^{2}+\varepsilon x_{1}x_{2}^{2}$ to define an isolated singularity.
2.2. Graph of a quasihomogeneous singularity
Let $f\in{\mathbb{C}}[x_{1},\dots,x_{N}]$ define an isolated singularity. Then for every index $j\leq N$ the polynomial $f$ has either the summand $x_{j}^{a}$ or a summand $x_{j}^{a}x_{k}$ for some exponent $a\geq 2$ and index $k\leq N$ (cf. [S71, Korollar 1.6] and [HK, Theorem 2.2]). Construct a map $\kappa:\{1,\dots,N\}\to\{1,\dots,N\}$. Set $\kappa(j):=j$ in the first case above and $\kappa(j):=k$ in the second.
Following [HK, Section 3] associate to $f$ the graph [^1] Such graphs were first considered by Arnold, however with the self-pointing arrows $j\to j$ too. We decide to remove such arrows to reduce complexity. $\Gamma_{f}$ with $N$ vertices labelled with the numbers $1,\dots,N$ and the oriented arrows $j\to\kappa(j)$ if $j\neq\kappa(j)$. In other words, the vertices correspond to the variables $x_{i}$ and the arrows to the monomials $x_{j}^{a}x_{k}$.
Example 5.

### 25 Mar 2024 v3

# Alexey BASALAEV ${}^{\rm ab}$ and Andrei IONOV ${}^{\rm c}$AuthorNameForHeadingA. Basalaev and A. Ionov

@@
[…]
Alexey BASALAEV ${}^{\rm ab}$ and Andrei IONOV ${}^{\rm c}$AuthorNameForHeadingA. Basalaev and A. Ionov
[…]
${}^{\rm a)}$ Faculty of Mathematics, National Research University Higher School of Economics,
[…]
${}^{\rm c)}$ Boston College, Department of Mathematics, Maloney Hall, Fifth Floor,
[…]
Received July 12, 2023, in final form March 06, 2024; Published online March 25, 2024
[…]
Consider the pairs $(f,G)$ with $f=f(x_{1},\dots,x_{N})$ being a polynomial defining a quasihomogeneous singularity and $G$ being a subgroup of ${\rm SL}(N,\mathbb{C})$, preserving $f$. In particular, $G$ is not necessary abelian. Assume further that $G$ contains the grading operator $j_{f}$ and $f$ satisfies the Calabi–Yau condition. We prove that the nonvanishing bigraded pieces of the B-model state space of $(f,G)$ form a diamond. We identify its topmost, bottommost, leftmost and rightmost entries as one-dimensional and show that this diamond enjoys the essential horizontal and vertical isomorphisms.
[…]
singularity theory; Landau–Ginzburg orbifolds
[…]
32S05; 14J33
1 Introduction
Let a polynomial $f\in{\mathbb{C}}[x_{1},\dots,x_{N}]$ be quasihomogeneous with respect to some positive integers ${d_{0},d_{1},\dots,d_{N}}$, i.e.,
$f\bigl{(}\lambda^{d_{1}}x_{1},\dots,\lambda^{d_{N}}x_{N}\bigr{)}=\lambda^{d_{0}}f(x_{1},\dots,x_{N}),\qquad\forall\lambda\in{\mathbb{C}}^{\ast}.$
Assume also that $x_{1}=\cdots=x_{N}=0$ is the only critical point of $f$ and $d_{1},\dots,d_{N}$ have no common factor. Then the zero set ${f(x_{1},\dots,x_{N})=0}$ defines a degree $d_{0}$ quasismooth hypersurface $X_{f}$ in ${\mathbb{P}}(d_{1},\dots,d_{N})$. Such hypersurfaces became of great interest in the early 90’s in the context of mirror symmetry (cf. [7, 8]). In particular, if the Calabi–Yau condition $d_{0}=\sum_{k=1}^{N}d_{k}$ holds, then the first Chern class of $X_{f}$ vanishes and, hence, its canonical bundle is trivial meaning that $X_{f}$ is a Calabi–Yau variety.
@@ 1 Introduction
1.1 Hodge diamonds of Calabi–Yau manifolds
The state space of a Calabi–Yau manifolds $X$ is the cohomology ring $H^{\ast}(X)$. This cohomology ring is naturally bigraded building up a Hodge diamond of size $D:=\dim_{\mathbb{C}}X$. In particular the following properties hold:
@@ 1.1 Hodge diamonds of Calabi–Yau manifolds
    there is a “vertical” vector space isomorphism $H^{p,q}(X)\cong\bigl{(}H^{D-q,D-p}(X)\bigr{)}^{\vee}$, where $(-)^{\vee}$ stands for the dual vector space
In physics, this bigrading is coming from considering Calabi–Yau B-model associated to $X$ and A-model bigrading is obtained from it by the so-called rotation of the diamond by $90^{\circ}$. On the level of state spaces the mirror symmetry map is an isomorphism of the cohomology vector spaces for a dual pair of Calabi–Yau manifolds switching A- and B-model bigradings.
More, generally if considering Calabi–Yau orbifolds in place of quasismooth varieties, one replaces the ordinary cohomology ring by the Chen–Ruan cohomology ring $H^{\ast}_{\rm orb}$. This is an essential question if $H^{\ast}_{\rm orb}$ forms a Hodge diamond too. Some of the Hodge diamond properties above follow directly from the definitions or could be verified in the similar way to our main Theorem 1.1 below, while the others (like the property (4)) were not investigated in literature up to our knowledge and do not look to be straightforward.
1.2 Landau–Ginzburg orbifolds
Another facet of mirror symmetry is given by matching the so-called Landau–Ginzburg orbifolds in place of Calabi–Yau manifolds or orbifolds (cf. [17, 22, 29, 30, 31]). Mathematically, these are the pairs $(f,G)$ with $f$ being a quasihomogeneous polynomial with the only critical point $0\in{\mathbb{C}}^{N}$ and $G$ being a group of symmetries of $f$.
@@ 1.2 Landau–Ginzburg orbifolds
${\mathrm{GL}}_{f}:=\left\{g\in{\mathrm{GL}}(N,{\mathbb{C}})\mid f(g\cdot{\bf x})=f({\bf x})\right\}.$
@@
$j_{f}\cdot(x_{1},\dots,x_{N}):=\bigl{(}{\rm e}^{2\pi\sqrt{-1}d_{1}/d_{0}}x_{1},\dots,{\rm e}^{2\pi\sqrt{-1}d_{N}/d_{0}}x_{N}\bigr{)}.$
@@
For any $G\subseteq{\mathrm{GL}}_{f}$, the pair $(f,G)$ is called a Landau–Ginzburg orbifold. One associates to it the state space, which is the $G$-equivariant generalization of a Jacobian ring of $f$, together with A- and B- model bigradings, which again differ by $90^{\circ}$-rotation from one another provided $G$ acts by transformations with determinant 1. Since these are interdependent, within this paper we will focus on just the B-model, as it has clearer geometric interpretation, much alike bigrading on the cohomology of Calabi–Yau manifold. For this reason, we will call this state space endowed with B-model bigrading by ${\mathcal{B}}(f,G)$.
Up till now, Landau–Ginzburg orbifolds were mostly investigated for the groups $G$ acting diagonally on ${\mathbb{C}}^{N}$ and also for $f$ belonging to a very special class of polynomials – the so-called invertible polynomials (cf. [4, 5, 6, 14, 19, 20, 21]). Also some work was done for the symmetry groups $G=S\ltimes G^{d}$ with $G^{d}$ acting diagonally and $S\subseteq S_{N}$ – a subgroup of a symmetric group on $N$ elements [2, 3, 10, 11, 12, 18, 24, 27, 32]. We relax both conditions in this paper.
Once again, the mirror symmetry attempts to match A- and B-state spaces for dual pairs of Landau–Ginzburg orbifolds. It should be mentioned, that the state space enjoys several other structures besides being just bigraded vector space, like multiplication or bilinear form on both A- and B-sides, which should also be compatible with the mirror map. They will not be considered in the present paper.
1.3 Calabi–Yau/Landau–Ginzburg correspondence
The final piece of matching comes from Calabi–Yau/Landau–Ginzburg correspondence, which relates respective A-models and their state spaces coming from Calabi–Yau and Landau–Ginzburg geometries of $(X_{f},G/J)$ and $(f,G)$ provided $J\subseteq G$. Mathematically, up till now this was proved for diagonal symmetry groups by Chiodo and Ruan [9, Theorem 14] and in some special cases with $N=5$ in [25].
@@ 1.3 Calabi–Yau/Landau–Ginzburg correspondence
Theorem 1.1.
Let $f\in{\mathbb{C}}[x_{1},\dots,x_{N}]$ be a quasihomogeneous polynomial satisfying Calabi–Yau condition and defining an isolated singularity. Then for any $G\subseteq{\mathrm{SL}}_{f}$, such that $J\subseteq G$ the state space ${\mathcal{B}}(f,G)$ forms a diamond of size $N-2$ in a sense that it satisfies conditions $(1)$$(6)$ above.
@@
The proof is summed up in Propositions 5.2, 5.3 and 5.4. The vertical and horizontal isomorphism are given in Section 4. ∎
2 Preliminaries and notations
2.1 Quasihomogeneous singularities
@@
$f\bigl{(}\lambda^{d_{1}}x_{1},\dots,\lambda^{d_{N}}x_{N}\bigr{)}=\lambda^{d_{0}}f(x_{1},\dots,x_{N}),\qquad\forall\lambda\in{\mathbb{C}}.$ (2.1)
@@
We will say that $f$ defines an isolated singularity at $0\in{\mathbb{C}}^{N}$ if 00 is the only critical point of $f$. According to K. Saito [28, Satz 1.3] one may consider without changing the singularity only the quasihomogeneous polynomials, such that $20<q_{k}\leq 1/$ 2 for all $k=1,\dots,N$. Moreover, we may assume that all variable $x_{k}$ with $q_{k}=1/2$ enter $f$ only in monomial $x_{k}^{2}$, in particular, there are no monomials of the form $x_{i}x_{j}$ with $i\neq j$. Then the number of its monomials is not less than $N$, the number of variables.
Example 2.1.
@@
$f$ $=x_{1}^{a_{1}}\qquad\text{Fermat type},$
$f$ $=x_{1}^{a_{1}}+x_{1}x_{2}^{a_{2}}+\dots+x_{N-1}x_{N}^{a_{N}}\qquad\text{chain type},$
$f$ $=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}x_{3}+\dots+x_{N-1}^{a_{N-1}}x_{N}+x_{N}^{a_{N}}x_{1}\qquad\text{loop type}.$
Using the word ‘type’, we mean the certain structure of the monomials without specifying the exponents $a_{i}$.
It’s easy to see that for Fermat, chain and loop type polynomials, the reduced weights $q_{1},\dots,q_{N}$ are defined in a unique way. This is also true for any quasihomogeneous singularity that we assume (cf. [28, Korollar 1.7]). We have
$\begin{split}&q_{1}=\frac{1}{a_{1}}\qquad\text{Fermat type,}\\ &q_{i}=\sum_{j=i}^{N}\frac{(-1)^{j-i}}{a_{1}\cdots a_{j}}\qquad\text{chain type,}\\ &q_{i}=(-1)^{N-1}\frac{1-a_{i}+a_{i}\sum_{k=2}^{N-1}(-1)^{k}\prod_{l=2}^{k}a_{i-l+1}}{\prod_{k=1}^{N}a_{k}-(-1)^{N}}\qquad\text{loop type,}\end{split}$
@@
Example 2.2 ([1, Section 13.1]).
If $N\leq 2$, then all quasihomogeneous isolated singularities are given by the $\oplus$-sums of the Fermat, chain or loop type polynomials.
Example 2.3 ([1, Section 13.2]).
If $N=3$, then all quasihomogeneous isolated singularities are given by the following polynomials $f_{\rm I}=x_{1}^{a_{1}}+x_{2}^{a_{2}}+x_{3}^{a_{3}}$, $f_{\rm II}=x_{1}^{a_{1}}+x_{2}^{a_{2}}x_{3}+x_{3}^{a_{3}}$, $f_{\rm III}=x_{1}^{a_{1}}+x_{2}^{a_{2}}x_{1}+x_{3}^{a_{3}}x_{1}+\varepsilon x_{2}^{p}x_{3}^{q}$, $f_{\rm IV}=x_{1}^{a_{1}}+x_{2}^{a_{2}}x_{3}+x_{3}^{a_{3}}x_{2}$, $f_{\rm V}=x_{1}^{a_{1}}+x_{2}^{a_{2}}x_{1}+x_{3}^{a_{3}}x_{2}$, $f_{\rm VI}=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}x_{1}+x_{3}^{a_{3}}x_{1}+\varepsilon x_{2}^{p}x_{3}^{q}$, $f_{\rm VII}=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}x_{3}+x_{3}^{a_{3}}x_{1}$ with some positive $a_{1}$, $a_{2}$, $a_{3}$. The numbers $a_{i}$ are arbitrary for $f_{\rm I}$, $f_{\rm II}$, $f_{\rm IV}$, $f_{\rm V}$, $f_{\rm VII}$, however the polynomials $f_{\rm III}$ and $f_{\rm VI}$ are only quasihomogeneous if $\varepsilon\neq 0$ and some additional combinatorial condition on $a_{1}$, $a_{2}$, $a_{3}$ holds. In particular the least common divisor of $(a_{2},a_{3})$ should be divisible by $a_{1}-1$ for $f_{\rm III}$ to exist. Allowed values of $\varepsilon$ depend quiet heavily on $a_{i}$. In particular $\varepsilon\not\in\{0,\sqrt{-1},-\sqrt{-1}\}$ for $f=x_{1}^{3}+x_{1}x_{2}^{2}+x_{1}x_{3}^{2}+\varepsilon x_{1}x_{2}^{2}$ to define an isolated singularity.
2.2 Graph of a quasihomogeneous singularity
Let $f\in{\mathbb{C}}[x_{1},\dots,x_{N}]$ define an isolated singularity. Then for every index $j\leq N$ the polynomial $f$ has either the summand $x_{j}^{a}$ or a summand $x_{j}^{a}x_{k}$ for some exponent $a\geq 2$ and index $k\leq N$ (cf. [28, Korollar 1.6] and [16, Theorem 2.2]). Construct a map $\kappa\colon\{1,\dots,N\}\to\{1,\dots,N\}$. Set $\kappa(j):=j$ in the first case above and $\kappa(j):=k$ in the second.
Following [16, Section 3] associate to $f$ the graph [^1] Such graphs were first considered by Arnold, however with the self-pointing arrows $j\to j$ too. We decide to remove such arrows to reduce complexity. $\Gamma_{f}$ with $N$ vertices labelled with the numbers $1,\dots,N$ and the oriented arrows $j\to\kappa(j)$ if $j\neq\kappa(j)$. In other words, the vertices correspond to the variables $x_{i}$ and the arrows to the monomials $x_{j}^{a}x_{k}$.
Example 2.4.